Problem V: Given v(x) = Sandw (x) = 10x + 3, find (vow)(x) and simplify.
Question 5
Answer #1

Need to find (vow)(x) & simplify it VCD & wWw= 10x+3 (vow)) V (w()) = V (10x+3) 6 1OX+3-3 (substitute x= 102tz in V(30) 10x 0